OpenStudy (anonymous):

Find the axis of symmetry of the graph of the function. f(x) = 2x^2 - 4x - 1

OpenStudy (anonymous):

\[-\frac{b}{2a}=-\frac{-4}{2\times 2}=1\] so \[x=1\] is your axis of symmetry

OpenStudy (anonymous):

thank you @satellite73